Twelve more than one-half the square of a number is 30. What is the number?
Answer by rm29924(97)

Let x = the number
make an equation
the square of a number is x^2
to get one-half of the square divide it by 2
x^2/2
twelve more means add 12
x^2/2+12=30
subtract 12 from both sides
x^2/2=18
to get rid of the 2 on the bottom multiply both sides by 2
x^2=36
to get x we have to take the square root of both sides
x=6
check your work
x^2+12=30
6^2/2+12=30
36/2+12=30
18+12=30
30=30